### v2

# Odd-Parity Altermagnetism Originated from Orbital Orders

Abstract
Odd-parity spin-splitting plays a central role in spintronics and unconventional superconductivity, yet its microscopic realization in collinear magnetic systems remains elusive. We propose a general symmetry-based strategy for realizing odd-parity altermagnetism by stacking two noncentrosymmetric monolayers in an interlayer antiferromagnetic configuration and applying an in-plane layer-flip operation. In this setting, odd-parity spin-splitting originates from nonrelativistic orbital orders rather than spin-orbit coupling, and is protected by an effective time-reversal symmetry despite the explicit time-reversal symmetry being broken. By exploiting lattice symmetries, our framework enables the realization of both $p$- and $f$-wave altermagnets. The resulting models generically host quantum spin Hall insulator phases, featuring topologically protected helical edge states and quantized spin Hall conductance. Our work expands the landscape of altermagnetic phases and opens a pathway toward spintronics and unconventional superconductivity in altermagnetic systems.
Momentum-dependent spin-splitting (MDSS) provides a fundamental link between the geometry of Bloch states and a diverse range of topological phases and spin transport phenomena. Depending on symmetry of the electron energy $E_{\boldsymbol{k},s}$, MDSS can be broadly classified into three classes [1, 2, 3, 4]: (i) even-parity, where $E_{\boldsymbol{k},s}=E_{-\boldsymbol{k},s}$, (ii) odd-parity, where $E_{\boldsymbol{k},s}=E_{-\boldsymbol{k},-s}$, and (iii) mixed-parity, where neither equality holds. Spin-orbit coupling (SOC) in noncentrosymmetric materials is a prototypical origin of MDSS [5, 6, 7]. As SOC is time-reversal even and inversion odd, the resulting spin-splitting always takes the odd-parity form. Over the past two decades, SOC-induced MDSS has stimulated extensive theoretical and experimental efforts [7, 8, 9, 10, 11, 12], including the discovery of various topological phases and exotic electromagnetic responses [13, 14, 15, 16, 17, 18].
Recently, a novel class of collinear antiferromagnets known as altermagnets [19, 20], has drawn wide interest for hosting nonrelativistic spin-split bands without net magnetization [20, 21, 22, 23, 24, 3, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and for their diverse associated phenomena [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. Within the context of spin group theory [20, 21, 62, 63, 64, 65], they are known to exhibit even-parity MDSS [20] (e.g., $d$-, $g$-, and $i$-wave), protected by the intrinsic symmetry $\bar{C}_{2}||\mathcal{T}]$ that imposes $E_{\boldsymbol{k},s}=E_{-\boldsymbol{k},s}$, where $\bar{C}_{2}$ is a $180^{\circ}$ rotation on spin about an axis perpendicular to it ($C_{2}$) combined with spin-space inversion (the overhead bar $\bar{\ \ }$”), and $\mathcal{T}$ denotes time-reversal in real space. In contrast, the realization of nonrelativistic odd-parity MDSS has received far less attention. So far, proposals mainly involve noncollinear antiferromagnets [66, 67, 68, 69], which are generally sensitive to external perturbations and therefore difficult to control. A more recent idea introduces odd-parity MDSS in bipartite collinear antiferromagnets [70], where sublattice currents generate phase-carrying hoppings akin to those in the Haldane model [71]. This breaks $\bar{C}_{2}||\mathcal{T}]$ but preserves $C_{2}||\mathcal{P}] [$72], with $\mathcal{P}$ the spatial inversion. Because $C_{2}||\mathcal{P}]$ enforces $E_{\boldsymbol{k},s}=E_{-\boldsymbol{k},-s}$, it leads to odd-parity MDSS in a collinear setting, referred to as odd-parity altermagnets [70]. Since their MDSS closely mimics the SOC-induced MDSS, these systems are expected to host a rich variety of interesting phenomena. Yet, because sublattice currents typically demand finely tuned flux patterns that are challenging to realize in real materials, developing a general and practical route to two-dimensional odd-parity altermagnets remains an open problem.
Electron hopping between atomic orbitals of different magnetic quantum numbers provides another mechanism to generate phase-dependent hopping [73, 6]. For example, in two-dimensional systems, when an electron hops from an orbital $a$ on one site to an orbital $b$ on another, the overlap integral can be expressed as
$t_{ab}(\boldsymbol{r}_{j})=\int\psi^{*}_{b}(\boldsymbol{r}-\boldsymbol{R}_{1})\mathcal{H}_{\rm hop}\psi_{a}(\boldsymbol{r}-\boldsymbol{R}_{2})d^{2}r,$ (1)
where $\psi_{a(b)}(\boldsymbol{r})$ are the orbital wave functions, $\boldsymbol{R}_{1(2)}$ are site positions, $\mathcal{H}_{\rm hop}$ is the hopping operator, and $\boldsymbol{r}_{j}=\boldsymbol{R}_{2}-\boldsymbol{R}_{1}$ is the bond vector. If the two orbitals have different magnetic quantum numbers $m_{a}$ and $m_{b}$, the hopping element along the bond $\boldsymbol{r}_{j}$ picks up a phase originating solely from the angular structure of the orbitals and depending on the azimuthal angle $\varphi_{j}$ of the bond direction. Specifically, one has $t_{ab}(\boldsymbol{r}_{j})=t_{j}e^{im\varphi_{j}}$, where $m=m_{a}-m_{b}$ and $t_{j}$ can be chosen as real. Crucially, this phase originates purely from nonrelativistic orbital order, rather than relativistic SOC. A well-known example is the Bernevig-Hughes-Zhang model for HgTe quantum wells [6, 74], in which the hopping between $s$- and $p$-orbitals yields effective SOC-like terms $A(\sin k_{x}\sigma_{x}+\sin k_{y}\sigma_{y})$ and produces topologically nontrivial band structures.
In this Letter, we present a bilayer stacking framework that realizes odd-parity altermagnetism driven by orbital order. The construction requires at least two orbitals with distinct magnetic quantum numbers and employs the symmetry $\bar{E}||\mathcal{M}_{z}\mathcal{T}]$ to connects the two layers, where $\mathcal{M}_{z}$ denotes mirror symmetry and $\bar{E}$ represents spin-space inversion via time-reversal [20, 75, 76, 77]. In two-dimensional layered systems, the symmetry $\bar{E}||\mathcal{M}_{z}\mathcal{T}]$ acts as an effective time-reversal symmetry and consequently guarantees $E_{\boldsymbol{k},s}=E_{-\boldsymbol{k},-s}$. Explicit tight-binding models show that this framework naturally yields both $p$- and $f$-wave altermagnets, and that the resulting systems can realize quantum spin Hall insulator phases characterized by helical edge states and quantized spin Hall conductance. This differs fundamentally from previous studies on systems with $C_{2}||\mathcal{P}]$, where one obtains Chern insulator phases characterized by chiral edge states.
General framework.–Generating odd-parity altermagnetism requires three key conditions. First, the system must exhibit a compensated collinear magnetic order [20]. Second, the bands must satisfy $E_{\boldsymbol{k},s}=E_{-\boldsymbol{k},-s}$, which can be guaranteed by an effective time-reversal symmetry or $C_{2}||\mathcal{P}]$. Third, the bands must satisfy $E_{\boldsymbol{k},s}\neq E_{-\boldsymbol{k},s}$ for generic momenta $\boldsymbol{k}$. In collinear antiferromagnets, the last condition implies the breaking of inversion symmetry within each spin sector, as well as the breaking of the intrinsic symmetry $\bar{C}_{2}||\mathcal{T}] [$20]. By contrast, in even-parity altermagnets, spin-splitting appears together with a preserved $\bar{C}_{2}||\mathcal{T}]$ and a typically conserved inversion symmetry.
To simultaneously fulfill the three conditions, here, we propose a bilayer stacking strategy, as illustrated in Fig. 1. We begin with two ferromagnetic monolayers stacked in an interlayer antiferromagnetic configuration. To satisfy the third condition, we consider noncentrosymmetric monolayers which lack inversion symmetry that enforces $E_{\boldsymbol{k},s}=E_{-\boldsymbol{k},s}$. In addition, at least two orbitals with different magnetic quantum numbers must be involved, such that phase-dependent hopping is present and $\bar{C}_{2}||\mathcal{T}]$ is explicitly broken. We then flip the top layer by an in-plane twofold rotation operation $C_{2i}$, where $i$ is aligned with a high-symmetry axis of the monolayer so that the lattice remains unchanged under the rotation. This operation reverses the orbital angular momentum $\mathcal{L}_{z}$, and therefore the magnetic quantum number $m$. The resulting top layer becomes a time-reversal copy of the bottom layer and the whole bilayer configuration preserves the symmetry $\bar{E}||\mathcal{M}_{z}\mathcal{T}]$, where $\mathcal{M}_{z}$ exchanges the two layers and $\mathcal{T}$ reverses the orbital angular momentum.
The above procedure naturally yields a band structure with odd-parity spin-splitting. Because the Fermi surface geometry is dictated by the lattice symmetry, designing the lattice structure provides a route to engineer various forms of odd-parity altermagnetism. For example, $p$- and $f$-wave altermagnets arise when the monolayer belongs to the $E||C_{1z}]$ and $E||C_{3z}]$ symmetry classes, respectively.
Figure 1: Illustration of the framwork. We begin with a noncentrosymmetric monolayer containing at least two orbitals with different orbital angular momenta $\mathcal{L}_{z}$ and $\mathcal{L}_{z}^{{}^{\prime}}$. Two such monolayers are stacked to form a bilayer exhibiting antiferromagnetic order. The top layer is then flipped by an in-plane $C_{2i}$ rotation, which reverses the orbital angular momentum while keeping the lattice unchanged. This procedure preserves the symmetry $\bar{E}||\mathcal{M}_{z}\mathcal{T}]$ but breaks both inversion and mirror symmetry $\mathcal{M}_{z}$ in real space, and thereby giving rise to odd-parity spin-splitting, as shown in the final stage.
Figure 2: Schematic of the lattice models for $p$- and $f$-wave altermagnets. Panels (a) and (e) show the noncentrosymmetric square and hexagonal lattices used for $p$- and $f$-wave altermagnets, respectively. (b) and (f) illustrate the phase-carrying inter-sublattice hoppings. The sign $\pm$ in the exponential refers to the top/ bottom layer. The resulting $p$- and $f$-wave patterns are displayed in (c-d) and (g-h), respectively. (c) [(g)] Spin-resolved Fermi surfaces at E=0.5E=-0.5 [$E=0.7$]. Common parameters for (c-d) [(g-h)] are $\{t_{1},t_{2},t_{0},t_{d},t_{\perp},\delta,J\}=\{1,0.6,0.2,0.2,0,0,1\}$. $\{t,t_{0},t_{d},t_{\perp},\delta,J\}=\{1,0.7,0.3,0,0,2\}$].
Models for $p$- and $f$-wave altermagnets.–We now implement the proposed framework in explicit lattice models. Guided by the symmetry requirements discussed above, the minimal setup consists of two layers, each containing two sublattices with distinct orbitals. This leads to an eight-band Hamiltonian of the form
$\mathcal{H}(\boldsymbol{k})$ $=$ $\left[\begin{array}[]{cc}\mathcal{H}_{t}(\boldsymbol{k})+J\boldsymbol{N}\cdot\boldsymbol{s}&t_{\perp}\\ t_{\perp}&\mathcal{H}_{b}(\boldsymbol{k})-J\boldsymbol{N}\cdot\boldsymbol{s}\\ \end{array}\right],$ (4)
where $\boldsymbol{s}=(s_{x},s_{y},s_{z})$ denotes the vector of Pauli matrices acting on spin space. For notational simplicity, we omit all identity matrices throughout the paper. The top and bottom layer Hamiltonians satisfy $\bar{E}||\mathcal{M}_{z}\mathcal{T}]\mathcal{H}_{t}(\boldsymbol{k})=\mathcal{H}_{b}^{*}(-\boldsymbol{k})$. The parameter $t_{\perp}$ describes the interlayer coupling, and $J$ is the strength of the interlayer antiferromagnetic exchange field.
To make the analysis transparent, we consider the simplest case where each sublattice in a monolayer hosts a pure orbital state. The generic Hamiltonian for top layer then becomes
$\mathcal{H}_{t}(\boldsymbol{k})$ $=$ $\sum_{j}t_{j}\left[{\rm Re}(\mathcal{F}_{j,\boldsymbol{k}})\sigma_{x}-{\rm Im}(\mathcal{F}_{j,\boldsymbol{k}})\sigma_{y}\right]$ (5)
$+\sum_{j}\left(t_{d}\sigma_{z}+t_{0}\right)\cos(\boldsymbol{k}\cdot\boldsymbol{b}_{j})+\delta\sigma_{z},$
where $\mathcal{F}_{j,\boldsymbol{k}}\equiv e^{i\boldsymbol{k}\cdot\boldsymbol{a}_{j}}e^{im\varphi_{j}}$ encodes the direction-dependent phase associated with the orbital angular momentum, and Pauli matrices $\sigma_{x,y,z}$ act on sublattices. The first row describes inter-sublattice hopping along the bond vectors $\boldsymbol{a}_{j}$, whereas the second row represents intra-sublattice hopping along $\boldsymbol{b}_{j}$ parameterized by $\{t_{0},t_{d}\}$ and the on-site potential difference $\delta$ between the two sublattices. We emphasize that the present model serves as a minimal realization of odd-parity altermagnetism driven by orbital order. Notably, a single monolayer with two sublattices hosting inequivalent orbitals and antiparallel local moments would generally become a compensated ferrimagnet [78, 79, 80] with finite net magnetization, because the distinct orbital character prevents the two sublattices being related by a simple symmetry. The bilayer configuration is therefore essential.
For $p$-wave altermagnets, the monolayer only needs to break inversion symmetry. A simple realization is provided by a noncentrosymmetric square lattice with sublattice dimerization. Here we assume that the dimerization is along the $x$-direction and that the two sublattices in the top (bottom) layer host $p_{+1(-1)}\equiv p_{x}\pm ip_{y}$ and $s$ orbitals, respectively [Fig. 2(a)], such that $|m|=1$. The bond vectors and their associated phases are listed in Table. 1. The resulting Hamiltonian can be expressed in terms of Pauli matrices as
$\mathcal{H}_{p}(\boldsymbol{k})$ $=$ $h_{1,\boldsymbol{k}}(t_{0}+t_{d}\sigma_{z})+h_{2,\boldsymbol{k}}\sigma_{x}+h_{3,\boldsymbol{k}}\sigma_{y}+\delta\sigma_{z}$ (6)
$+(h_{4,\boldsymbol{k}}\sigma_{y}+h_{5,\boldsymbol{k}}\sigma_{x}+J\boldsymbol{N}\cdot\boldsymbol{s})\rho_{z}+t_{\perp}\rho_{x},$
where the Pauli matrices $\rho_{x,y,z}$ act on the two layers. The explicit forms of the functions $h_{i,\boldsymbol{k}}$ are summarized in Table. 2. In this configuration, the combined mirror operation $\mathcal{M}_{y}\mathcal{M}_{z}$ relates the two layers and enforces spin degeneracies along $\mathcal{M}_{y}$-invariant lines in the Brillouin zone [Fig. 2(c,d)]. By contrast, if the dimerization is oriented along $y$, the relevant symmetry becomes $\mathcal{M}_{x}\mathcal{M}_{z}$, resulting in a $\pi/2$ rotation of the spin-split pattern [see Supplemental Materials (SM) [81]]. Hence, the orientation of the monolayer dimerization provides an efficient knob to control the $p$-wave spin-split pattern.
Bond vectors Complex phases $\varphi_{i}$
Square
$\boldsymbol{a}_{1}=a\{1,0\}$, $\boldsymbol{a}_{2}=a\{0,1\}$,
$\boldsymbol{a}_{3}=a\{-1,0\}$, $\boldsymbol{a}_{4}=a\{0,-1\}$
$(i-1)\pi/2$
$\boldsymbol{b}_{1}=\boldsymbol{a}_{1}+\boldsymbol{a}_{2}$, $\boldsymbol{b}_{2}=\boldsymbol{a}_{2}+\boldsymbol{a}_{3}$,
$\boldsymbol{b}_{3}=\boldsymbol{a}_{3}+\boldsymbol{a}_{4}$, $\boldsymbol{b}_{4}=\boldsymbol{a}_{4}+\boldsymbol{a}_{1}$
0
Hexagonal
$\boldsymbol{a}_{1}=a\{0,1\}$, $\boldsymbol{a}_{2}=a\{\frac{\sqrt{3}}{2},-\frac{1}{2}\}$,
$\boldsymbol{a}_{3}=a\{-\frac{\sqrt{3}}{2},-\frac{1}{2}\}$
$(i-1)2\pi/3$
$\boldsymbol{b}_{1}=\boldsymbol{a}_{2}-\boldsymbol{a}_{3}$, $\boldsymbol{b}_{2}=\boldsymbol{a}_{3}-\boldsymbol{a}_{1}$,
$\boldsymbol{b}_{3}=\boldsymbol{a}_{1}-\boldsymbol{a}_{2}$
0
Table 1: Bond vectors for inter(intra)-sublattice hopping $\boldsymbol{a}_{i}$ ($\boldsymbol{b}_{i}$), and the associated complex phases for square and hexagonal lattices shown in Fig. 2. The lattice constant $a$ is set to unity throughout.
$p$-wave $f$-wave
$h_{1,\boldsymbol{k}}$ $4\cos k_{x}\cos k_{y}$ $\cos\sqrt{3}k_{x}+2\cos\frac{\sqrt{3}k_{x}}{2}\cos\frac{3k_{y}}{2}$
$h_{2,\boldsymbol{k}}$ $(t_{1}-t_{2})\cos k_{x}$ $t(\cos k_{y}-\cos\frac{\sqrt{3}k_{x}}{2}\cos\frac{k_{y}}{2})$
$h_{3,\boldsymbol{k}}$ $(t_{1}+t_{2})\sin k_{x}$ $t(\sin k_{y}+\cos\frac{\sqrt{3}k_{x}}{2}\sin\frac{k_{y}}{2})$
$h_{4,\boldsymbol{k}}$ 0 $\sqrt{3}t\sin\frac{\sqrt{3}k_{x}}{2}\sin\frac{k_{y}}{2}$
$h_{5,\boldsymbol{k}}$ $2t_{y}\sin k_{y}$ $\sqrt{3}t\sin\frac{\sqrt{3}k_{x}}{2}\cos\frac{k_{y}}{2}$
Table 2: The functions $h_{i,\boldsymbol{k}}$ for $p$- and $f$-wave altermagnets.
The $f$-wave case additionally requires that the monolayer respects $E||C_{3z}]$ symmetry. To this end, we consider a hexagonal lattice in which the two sublattices are occupied $p_{+1(-1)}$ and $s$ orbitals in the top (bottom) layer, respectively [Figs. 2(e-f)]. The resulting Hamiltonian has the same structure as Eq. (6), but with different expressions for the functions $h_{i,\boldsymbol{k}}$ (see Table. 2). As illustrated in Fig. 2(g,h), two $C_{3}$-symmetric spin-polarized Fermi surfaces are related by a $\mathcal{M}_{x}$ operation, giving rise to an $f$-wave spin-split pattern. In contrast to the $p$-wave case, the $C_{3}$ symmetry forces the Fermi surfaces to enclose at least one of the three $C_{3}$-invariant momenta, i.e., $\boldsymbol{\Gamma}$, K, and $\textbf{K}^{\prime}$. Consequently, by tuning the chemical potential, the Fermi surface can be shifted from $\boldsymbol{\Gamma}$ to $\textbf{K}/\textbf{K}^{\prime}$, opening up the possibility of valley-polarized spin transport [82, 83, 84].
Figure 3: (a) The $\tau\rho=1$ sectors host two spin-polarized massive Dirac cones separated by a gap $\Delta E=2J$ at each valley. The dashed (solid) lines refer to bands with a Chern number of $C_{s,\tau}=-1/2$ ($C_{s,\tau}=1/2$); the color represents the spin polarization. (b) The spin-resolved Fermi surfaces are obtained at Fermi energy $\mu_{1}=-1.5t_{0}+J+1.4$ and $\mu_{2}=-1.5t_{0}-J+1.4$ when $\{t_{0},t_{\perp},J\}=\{0.1,0,2\}$. (c) Mid-gap helical edge states. $\{t_{0},t_{\perp},J\}=\{0.1,0.1,2\}$. (d) Topological phase diagram as a function of interlayer coupling $t_{\perp}$ and exchange field $J$. $t_{0}=0$. The spin Chern number becomes $C^{s}=+1$ (1-1) when the chemical potential lying in the bulk energy gap between the two upper (lower) bands. Common parameters for (b-d): $\{t,t_{d},\delta\}=\{1,0.5,0\}$.
Altermagnetic quantum spin Hall insulators.–In our framework, both the $p$- and $f$-wave models can realize quantum spin Hall insulator phases. The intuition is clear in the limit of vanishing interlayer coupling: the two layers decouple and behave as independent ferromagnetic monolayers that break time-reversal symmetry. When parameters fall into the topological regime, each monolayer becomes a Chern insulator. Because the two layers are related by an effective time-reversal symmetry, their Chern numbers cancel, leving a zero total (charge) Chern number but a finite spin Chern number [85], thus realizing a quantum spin Hall insulator.
To make the discussion concrete, we focus on the $f$-wave model (the $p$-wave case is presented in SM [81]). Setting $t_{\perp}=0$ makes the layer index $\rho_{z}$ a good quantum number. Expanding the Hamiltonian around the valleys K and $\textbf{K}^{\prime}$ up to linear order in momentum $\boldsymbol{q}$, we obtain
$\mathcal{H}_{f,\tau}^{s,\rho_{z}}(\boldsymbol{q})$ $=$ $s\rho J-\frac{3t_{0}}{2}+\left(\delta-\frac{3t_{d}}{2}\right)\sigma_{z}+\frac{3t(1-\tau\rho)}{2}\sigma_{x}$ (7)
$+\frac{3t(1+\tau\rho)}{4}(\tau q_{x}\sigma_{x}+q_{y}\sigma_{y}),$
where $\tau=\pm 1$ labels K and $\textbf{K}^{\prime}$ valleys, and $\rho=\rho_{z}=\pm 1$ denotes the top and bottom layers. The $\tau\rho=-1$ sectors are topologically trivial due to the vanishing linear dispersion. In contrast, the $\tau\rho=1$ sectors host two spin-polarized massive Dirac cones separated by a gap $\Delta E=2J$, as schematically shown in Fig. 3(a). Each Dirac cone carries a Chern number $\mathcal{C}_{s,\tau}=-{\rm sgn}(\tau t_{d})/2$ for $\delta<3|t_{d}|/2$. Because the effective time-reversal symmetry $\bar{E}||\mathcal{M}_{z}\mathcal{T}]$ relates K valley in the top (bottom) layer to the $\textbf{K}^{\prime}$ valley in the bottom (top) layer, Dirac cones with opposite spin and layer indices appear at the same energy. The system therefore becomes a quantum spin Hall insulator, characterized by the spin Chern number $\mathcal{C}^{s}=\frac{\mathcal{C}_{\uparrow,\rho_{z}}-\mathcal{C}_{\downarrow,-\rho_{z}}}{2}$, and hosts a pair of helical edge states within each of the two bulk gaps located at $E=-\frac{3t_{0}}{2}\pm J$, as shown in Fig. 3(c).
Importantly, this topological character survives finite interlayer coupling $t_{\perp}$, provided that the two upper (lower) bands in each spin sector are separated by the gaps around $E=-\frac{3t_{0}}{2}\pm J$ [see Fig. 3(d)]. Although $t_{\perp}$ hybridizes the layers and removes $\rho_{z}$ as a good quantum number, the spin Chern number remains unchanged as long as these gaps stay open. In practice, this situation arises when the exchange field $J$ is comparable to, or larger than, the band width $W$ at $t_{0}=0$. For $J\sim W$, a small $t_{\perp}$ can open the band gaps that are closed at $t_{\perp}=0$, driving the system into a topological phase, while for larger $J$ the gap is already stablized by the exchange field. Once the system is nontrivial, further increasing $t_{\perp}$ eventually closes and reopens the gaps, rendering the system trivial. We emphasize that this topologically nontrivial behavior is a generic feature of odd-parity altermagnets in our framework, and stands in clear contrast to previously proposed noncollinear $p$-wave magnets [66, 67], which are typically trivial.
Figure 4: Spin-split patterns and zero-temperature spin Hall conductance under different degree of orbital mixing. (a) Comparison of the spin-split patterns for $\eta=1$ and $\eta=-1$ at the same energy (${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}E}=-1.5t_{0}+J+1.4$). (b) Spin Hall conductance (in units of $e/2\pi$) as a function of chemical potential $\mu$, showing a sign reversal as $\eta$ across $\eta_{c}=0$. The parameters are chosen as $t+t^{\prime}=1$, $t_{0}=t_{\perp}=0.1$, $t_{d}=0.5$, $\delta=0$, and $J=2$.
Effects of orbital mixing.–In realistic systems, the electronic states at a given lattice site are typically linear combinations of different atomic orbitals. Since the inter-orbital hoppings are essential for generating odd-parity MDSS and nontrivial topology, it is important to assess how orbital mixing influences the altermagnetic phase. When orbital mixing is present, all hopping processes between the occupied orbitals of different sublattices must in principle be taken into account. Nevertheless, we find that only a subset of these hoppings actually matters, and that this subset is dictated purely by lattice symmetry. For example, on a square lattice, the phase associated with a hopping is $e^{im(j-1)\frac{\pi}{2}}$, which becomes trivial when $m=2n$ with $n\in\mathbb{Z}$ (since $e^{in\pi}=\pm 1$). Therefore, only hoppings between orbitals with odd angular momentum difference ($m\in 2\mathbb{Z}+1$) contribute to the odd-parity MDSS. A similar argument holds in systems with $C_{3}$ symmetry, where the associated phase becomes trivial for $m\in 3\mathbb{Z}$, and only hoppings with $m\notin 3\mathbb{Z}$ are relevant.
To illustrate this more concretely, let us revisit the $f$-wave altermagnet shown in Fig. 2(e). Suppose the purple sublattices in the top (bottom) layer now host a mixture of $s$ and $p_{-}$ ($p_{+}$) orbitals. This introduces additional hopping terms between $p_{-}$ and $p_{+}$ orbitals,
$\mathcal{H}^{\prime}_{f}(\boldsymbol{k})$ $=$ $f_{2,\boldsymbol{k}}^{\prime}\sigma_{x}+f_{3,\boldsymbol{k}}^{\prime}\sigma_{y}-f_{4,\boldsymbol{k}}^{\prime}\sigma_{y}\rho_{z}-f_{5,\boldsymbol{k}}^{\prime}\sigma_{x}\rho_{z},$ (8)
where $f_{i,\boldsymbol{k}}^{\prime}=\frac{(1-\eta)t}{1+\eta}f_{i,\boldsymbol{k}}$ ($i=2,3,4,5$), and $\eta\equiv\frac{t-t^{\prime}}{t+t^{\prime}}$ measures the strength of orbital mixing. Here $t^{\prime}$ refers to the amplitude of the $p_{-}\leftrightarrow p_{+}$ hopping. As $\eta$ is reduced from $\eta=1$, the spin-splitting gradually decreases and eventually vanishes at the critical value $\eta_{c}=0$, where an emergent $C_{2}||\mathcal{M}_{z}]$ symmetry suppresses the $f$-wave spin-splitting entirely [86, 87]. Remarkably, upon further decreasing $\eta$ beyond $\eta_{c}$, the spin-splitting reappears with a reversed pattern, corresponding to a time-reversed partner of the $\eta>0$ regime [see Fig. 4(a)]. It associates with a change of topology. As a direct consequence, the spin Hall conductance [8] reverses its sign across the transition [81], as shown in Fig. 4(b).
Discussions and conclusions.—Using symmetry principles as a guide, we demonstrate that odd-parity altermagnetism can emerge from the interplay of orbital order and collinear antiferromagnetism. This mechanism yields quantum spin Hall insulator phases when the exchange field becomes comparable to the band width, a regime achievable in magnetic van der Waals (MvdW) compounds. In our construction, because the layer-flip operation does not depend on specific monolayer symmetries, any MvdW system with interlayer antiferromagnetism [88, 89, 90, 91, 92, 93] can in principle be converted to an odd-parity altermagnet via symmetry engineering. Magnetic Janus monolayers are a representative family of materials that naturally fulfill the required symmetry conditions, as their structure inherently lacks both inversion and mirror symmetries [94]. The defining characteristic of odd-parity altermagnets, i.e., odd-parity spin-splitting, can be directly detected in experiments using spin-resolved ARPES [28, 27, 29, 95].
Beyond their fundamental interest, odd-parity altermagnets also provide a promising platform for unconventional superconductivity. Although magnetism inherently breaks time-reversal symmetry, the Fermi surfaces still obey the relation $E_{\boldsymbol{k},s}=E_{-\boldsymbol{k},-s}$. This condition is compatible with spin-singlet Cooper pairing, making the coexistence of superconductivity and magnetism a distinct possibility in these systems. Notably, recent studies have shown that in noncollinear $p$-wave magnets [66, 67], the interplay between $p$-wave spin-splitting and superconductivity leads to exotic phenomena [96, 97, 98, 99, 100, 101, 102, 103, 104]. Given that odd-parity altermagnets host similar Fermi-surface structures but within a more stable collinear magnetic environment, they provide an ideal and potentially superior playground for exploring similar and even richer superconducting physics.
In summary, we have developed a universal and experimentally feasible framework for realizing odd-parity altermagnetism. By extending the concept of altermagnets from the even-parity to the odd-parity regime, our work significantly broadens the landscape of altermagnetic materials and opens new avenues for research in spintronics and unconventional superconductivity.
Note added.–During the preparation of this paper, we became aware of other independent works [105, 106, 107, 108] that discuss the generation of odd-parity spin-splitting in 2D and 3D collinear antiferromagnets. Their approach relies on Floquet engineering to generate Haldane-like phase-dependent hopping via the application of time-periodic light. The odd-parity spin-splitting in their setting is all enforced by the $C_{2}||\mathcal{P}]$ symmetry, rather than the $\bar{E}||\mathcal{M}_{z}\mathcal{T}]$ symmetry emphasized in our framework.
Acknowledgements.— Z.-Y.Z. would like to thank Song-Bo Zhang and Luigi Camerano for helpful discussions, and the organizers of the “Workshop on altermagnets” for providing valuable opportunities for academic exchange. Z.-Y.Z., D.Z., D.L., and Z.Y. are supported by the National Natural Science Foundation of China (Grant No. 12174455), and Guangdong Basic and Applied Basic Research Foundation (Grant No. 2023B1515040023). Z. W. is supported by the National Natural Science Foundation of China (Grant No. 12474264), Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2404007), National Key R&D Program of China (Grant No. 2022YFA1404103).